a train starting from rest is accelerated and the instantaneous acceleration is given by 10/v+1m/s^2 where v is the velocity in m/s find the time after which the body attains a velocity of 54 km/hr

Grade:11

4 Answers

Arun
5 years ago
Dear student
 
a = 10/v+1 = dv/dt
10 * dt = (v+1) dv
On integrating
10 t = v² /2 + v + c
At t = 0 , v = 0 hence c= 0
Now
10 t = 15²/2 + 15
t = 112.5 +15 /10 = 12.75 sec
Now you can find the distance
You will get the distance as 38.33 m
 
Regards
Arun (askIITians forum expert)
Khimraj
5 years ago
a = 10/(v + 1)
dv/dt = 10/(v + 1)
(v+1)dv = 10dt
v2/2 + v =10t
v = 54km/h = 15m/s
So
225/2 + 15 = 10t
255/2 = 10t
255/20 = t
t = 12.75 sec.
 
Koushal malakar
5 years ago
A train starting from rest is accelerated and the instantaneous acceleration is given by 10/v+1 m/s2 .where v is the velocity in m/s. Find the distance in which the train attains a velocity 54 km/hr.
 
Koushal malakar
5 years ago
A train starting from rest is accelerated and the instantaneous acceleration is given by 10/v+1 m/s2 .where v is the velocity in m/s.Find the distance in which the train attains a velocity 54 km/hr.